Please help me solve this ratio word problem.
The lengths of the side of a triangle are in the extended ratio 3:10:12.
The perimeter is 400 cm. What is the length of the longest side in centimeters?

Answer by checkley79(3341)

extended ratio 3:10:12
P=A+B+C
3+10+12=25
400/25=16 IS THE MULTIPLIER OF THE RATIOS.
12*16=192 CM. IS THE LONGEST SIDE.
PROOF:
400=3*16+10*16+12*16
400=48+160+192
400=400